# A Theory of Recommendations

### v2

@@
A Theory of Recommendations
Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch. Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch. We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
(October 11, 2024)
This paper investigates the value of recommendations for disseminating economic information, with a focus on frictions resulting from preference heterogeneity. We consider Bayesian expected-payoff maximizers who receive non-strategic recommendations by other consumers. The paper provides conditions under which different consumer types accept these recommendations. Moreover, we assess the overall value of a recommendation system and the determinants of that value. Our analysis highlights the importance of disentangling objective information from subjective preferences when designing value-maximizing recommendation systems.
@@ 1 Introduction
Recommendations play an essential role for the diffusion of economic information. Job applicants frequently rely on assessments of potential employers by their friends. Conversely, job candidates often have to provide letters of recommendation from former employers. Consumers may base their choices of travel destinations, restaurants, or doctors on what they have heard from their acquaintances. In the digital economy, product ratings, typically reflecting anonymous consumer reports, are omnipresent.
Despite their relevance, recommendations are unlikely to function smoothly. For instance, too few recommendations may be given (Kremer et al., 2014; Che and Hörner, 2018), recommenders may be positively selected (Acemoglu et al., 2022), and interested parties might interfere with the process by choosing which recommendations to publish (Bolton et al., 2013; Tadelis, 2016). In the literature, it is common to make the simplifying assumption that product qualities can be objectively ranked. [^1] See also Che and Hörner (2018) and Hui et al. (2023). Instead, our paper assumes that preferences are heterogeneous, in which case product qualities cannot be objectively ranked and recommendations may be hard to interpret, even without strategic biases. We ask whether and how recommendations can be useful in spite of preference heterogeneity. Even well-intended recommendations by non-strategic players may reflect the senders’ preferences rather than the objective truth. A receiver who does not question the informational content of a recommendation may therefore make biased decisions.
To motivate our approach, consider a consumer’s choice between medical doctors. This is a complex problem, as the perceived quality of medical care will typically depend on many aspects. How carefully does a doctor listen to the patient? Is the technical equipment up to date? Is the staff friendly? How crowded is the waiting room? Moreover, different people may weigh different aspects differently. Despite such preference heterogeneity, rather than randomly choosing a doctor, people often base their decision on recommendations. Importantly, senders have limited stakes in the process and therefore lack incentives to think through the effects of their recommendation carefully. By contrast, adequately using the information can be crucial for the receiver. Finally, even if the sender is willing to think through the recommendation carefully, it will typically not be possible to convey all relevant information. For all these reasons, it is not obvious why or when recommendation can be useful.
This example illustrates three key features of our theory of recommendations:
Heterogeneity of Products and Preferences: Products come in different versions and consumers value them differently.
Asymmetry of Stakes: Making sure that the recommendation is not misunderstood is more important for the receiver than for the sender.
Coarseness of Recommendations: The sender is unable to provide enough detail for all relevant aspects of the service to become fully transparent to the receiver.
These features will be even more salient for ratings provided on internet platforms, as such ratings usually do not come from close acquaintances. Hence, the receivers should consider that senders have preferences which differ from their own and are likely to give an incomplete account of their consumption experience without exerting much effort. Backing out the implications of the recommendation for the receiver’s own decision is therefore a complex task. Relevant considerations for the interpretation are these: To what extent does the recommendation contain objective truth? And when it does not, how likely is it that the preferences of the sender and receiver are aligned? The answers to these questions will determine whether it is in the receiver’s interest to accept the recommendation and how valuable it is.
@@
Our model captures the above three features of recommendations in the simplest possible way. An agent must decide between several options, referred to as products for ease of exposition. Products are two-dimensional objects, differing only in whether they have high or low quality in each dimension. Thus, four product versions exist. There is a continuum of agent types. All types agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good. Conversely, they agree that a product is not worth choosing if it has low quality in both dimensions and is thus (objectively) bad. Accordingly, we call the products uncontroversial in both of these cases. By contrast, we refer to products as controversial when they have high quality in one dimension and low quality in the other. An agent prefers a controversial product that has high quality in the dimension that is more relevant for them, and is thus subjectively good, to a product that has high quality only in the less-relevant dimension (thus subjectively bad). Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences (as captured by the continuous type distribution).
@@
We assume that products are experience goods so that each agent only has stochastic knowledge of the quality distribution before consumption: They know the probability of each product version. The agent obtains a recommendation from a previous consumer of one of the products. The recommendation reveals whether the sender had a consumption experience that was sufficiently good (a buy recommendation), with payoffs above an exogenous threshold level $R$, or not (a don’t-buy recommendation). Independent of $R$, the sender (she) always gives a buy recommendation for a good product and a don’t-buy recommendation for a bad product. In addition, she sends a buy recommendation for a controversial product if her preferences for the product are strong enough. In line with her low stakes in the issue, the sender thus only provides a coarse description of her own consumption experience. By contrast, the receiver (he) of the recommendation, with his high stakes in the issue, will carefully evaluate the informational content of the recommendation using Bayesian inference, taking into account the distribution of product qualities, his own preferences, and those of the population.
In this framework, we address a variety of questions: How do recommendations create value for receivers, how can we interpret this value, and when is it maximal? If one could influence the payoff threshold beyond which a sender gives a buy recommendation, how high should it be? Are recommendations valuable even when uncontroversial products are unlikely? How does the value depend on how polarized the population is?
To answer these questions, we need to derive intermediate results of independent interest. Most importantly, we investigate under which conditions receivers accept all recommendations. We find that for a Bayesian expected-payoff maximizer, two key quantities measuring the objective and subjective content of the recommendation are decisive for receiver behavior. These quantities are simple functions of the model primitives. Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be (objectively) good and less likely to be bad. The subjective content captures the effect of the recommendation on the relative probability of the two controversial product versions. The evaluation of the subjective content thus depends on the receiver’s type. When the recommendation bears little or no subjective content, all types accept all recommendations, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation. As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept all recommendations. When the recommendation threshold $R$ approaches the maximal payoff, a buy recommendation essentially becomes objective, and similarly for a don’t-buy recommendation when $R$ approaches the minimal payoff. As receivers are aware of this, they accept all recommendations in each of these polar cases since the subjective component vanishes.
Based on this analysis, we then study the value of recommendation systems. We think of this value as the expected-payoff increase of a randomly chosen receiver resulting from a recommendation given by a randomly chosen sender. When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept all recommendations, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all. Therefore, only buy recommendations create value—they do so by inducing the receiver to purchase the product if it generates a higher expected payoff for him than the outside option. When the recommendation contains more subjective content, only a subset of the receivers accept it, rendering the analysis more complex. The value of the recommendation system then has two sources. First, receivers who accept all recommendations benefit because they expect the recommended product to be better than the outside option. Second, receivers who never accept recommendations benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
Next, we turn to the design of recommendation systems. Motivated by the practice of digital platforms, we consider the recommendation threshold $R$ as a design variable. When the type distribution is symmetric, so that there is no preference bias on average, the value of the recommendation system is a monotone function of the threshold. If the value is increasing, the product optimally receives a buy recommendation only if it is good; if the value is decreasing, the product optimally receives a don’t-buy recommendation only if it is bad. The former case arises if good products are more likely ex ante than bad products are, and conversely for the latter case. Either way, the content of the recommendation essentially becomes objective. In contrast, with an asymmetrically distributed population, value maximization does not necessarily require an extreme threshold. Rather than revealing only whether a product is good or bad, it may be optimal for the sender to give a buy recommendation even if the product is controversial but sufficiently aligned with her preferences. Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common and there is broad agreement on which of them is preferable. By contrast, extreme recommendation thresholds that only reveal objective content are still desirable if the product is much more likely to be objectively good than objectively bad or vice versa.
We then discuss further determinants of the value of the recommendation system, assuming that the population is symmetric. First, we find that an increase in the probability of controversial products lowers the value only if the probabilities of the two controversial versions are similar. Second, whether a mean-preserving spread in the population distribution (corresponding to a more polarized population) increases the value of a given recommendation system depends on how strict the recommendation threshold is and on how common good products are relative to bad ones.
Finally, we show the robustness of our results. With mild adjustments, the main insights continue to hold when sender and receiver distributions are distinct—a situation that could arise because, for instance, recommendations stem from early adopters who do not represent the population at large. This also has non-trivial design implications: Our results on the relation between mean-preserving spreads can be adapted to show how a designer would want to choose the sender distribution, which may be at least partly under his control. Second, we allow for more than two types of recommendations. Again, extreme recommendation thresholds increase the probability that a recommendation has objective content and thus increase the value of the system when the population is symmetrically distributed. Third, we show that having access to many recommendations does not necessarily change receiver behavior compared to the case of a single recommendation with an optimal threshold. Then, having access to multiple recommendations does not increase the value.
Section 2 relates the paper to the literature. Section 3 introduces the model and discusses its key assumptions. In Section 4, we describe the conditions under which different receiver types accept all recommendations. Then, in Section 5 we first characterize the value of a recommendation system before investigating the determinants of the value and the design of value-maximizing systems. Section 6 discusses our results and considers extensions. Section 7 concludes. All proofs are relegated to the appendix.
2 Relation to the Literature
Our paper investigates frictions in recommendation systems arising from preference heterogeneity, emphasizing the behavior of receivers. By contrast, the literature has mainly focused on frictions on the sender side.
Public Goods Problems
@@ 2 Relation to the Literature
Research on recommendation systems often deals with the exploration-exploitation dilemma and its implications for providing recommendations, interpreted as the provision of a public good. Several authors argue that consumers usually neglect the positive externalities generated by their reviews, leading to insufficient exploration and reviewing, in particular for lesser-known niche products (e.g., Kremer et al. (2014) and Che and Hörner (2018)). Expanding on these papers, Vellodi (2022) considers the impact on market structure, showing how unbiased reviews might disadvantage entrants due to the “cold start” problem, thereby reinforcing the monopoly position of incumbents. These studies view the design of the recommendation system as the solution of a public goods problem by appropriate information management. [^2] For example, Kremer et al. (2014) and Che and Hörner (2018) suggest reducing transparency or endorsing yet-to-be-proven products, while Vellodi (2022) supports the idea of withholding reviews from incumbents to encourage new participants.
Sender Selection
The selection of recommenders is another source of frictions in recommendation systems. In their analysis of the learning dynamics resulting from recommendations, Acemoglu et al. (2022) argue that consumers with strong initial preferences for a product are more likely to purchase it. [^3] In contrast to our paper, their question is essentially whether and how fast consumers find out if the product is good or not, an assessment with which all consumers agree. Beyond that, we model preference heterogeneity differently. Recommendations thus inevitably carry an inherent bias reflecting the predispositions of the recommender pool. [^4] In Section 6.1, we allow for distinct sender and receiver populations and thus for the possibility of a bias among the recommenders relative to the overall population. Receivers should recognize this bias rather than accepting recommendations at face value. Our paper intersects with theirs by highlighting the significance of preference heterogeneity for interpreting recommendations. However, contrary to Acemoglu et al. (2022), we focus on the determinants of the value of recommendation systems and the design of those systems rather than the possibility and speed of learning. [^5] In Section 6.3, we introduce multiple recommendations, as do Acemoglu et al. (2022), but, in contrast to them, we ask under which conditions multiple recommendations enhance system value.
Non-truthful Recommendations and Strategic Manipulations
@@
Strategic actions of recommenders, firms, and platforms may further reduce the value of recommendations for consumers. Bolton et al. (2013) have shown that eBay users often refrain from negative feedback to avoid retaliation. [^6] Tadelis (2016) discusses the design implications of these issues—using the example of eBay deciding to refrain from allowing seller evaluations of buyers—and also touches on strategic motives in recommendations more generally. Like us, Chakraborty and Harbaugh (2010) assume that the sender possesses multidimensional information. They show how this helps her influence the behavior of the receiver in spite of a common bias. Instead, we emphasize the effects of preference heterogeneity without strategic behavior, assuming that senders mechanically provide truthful recommendations. In contrast with our model with consumer recommendations, in Peitz and Sobolev (forthcoming) and Johnen and Ng (2024) firms engage in sophisticated strategies to inflate their ratings. [^7] Peitz and Sobolev (forthcoming) investigate firm-driven recommendations and the choice of sales channel: targeted (with recommendation) or non-targeted. They identify motives behind “inflated recommendations,” where the firm also recommends welfare-reducing bad product matches. Johnen and Ng (2024) analyze a related phenomenon, where firms initially lower prices to attract positive ratings, later leveraging these ratings to justify price increases—effectively reducing the reliability of ratings as a quality indicator. More closely related to our paper, Bourreau and Gaudin (2022) consider incentives for firms to manipulate recommendations in a setting with heterogeneous preferences.

### v3

# A Theory of “Likes”

@@
A Theory of “Likes”
Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch. Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch. This paper was previously circulated under the title “A Theory of Recommendations”. We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Matt Mitchell, Marc Möller, Nick Netzer, Volker Nocke, Marcus Reisinger, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at MaCCI (2025), SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
(May 1, 2025)
This paper investigates the value of recommendations for disseminating economic information, with a focus on frictions resulting from preference heterogeneity. We consider Bayesian expected-payoff maximizers who receive non-strategic recommendations by other consumers. We show by which channels a recommendation system influences consumer demand and welfare. Our analysis reveals how the welfare-maximizing design of a recommendation system may differ from what a profit-maximizing designer would choose to do.
@@ 1 Introduction
Recommendations play an essential role for the diffusion of economic information in diverse domains. For instance, consumers may base their choices of travel destinations, restaurants, or doctors on what they have heard from their acquaintances. In the digital economy, the supply of recommendations is increasingly mediated through simple mechanisms such as “likes,” star ratings, or thumbs up/down. These coarse signals compress personal evaluations into binary or low-dimensional feedback that guides others’ decisions.
In this paper, we ask how recommendations can function when preferences are heterogeneous and product qualities cannot be objectively ranked: is a “like” a reflection of inherent quality or merely an expression of individual taste? Consider a user of a streaming platform deciding whether to “like” a documentary. Perceived quality depends on many aspects—informativeness, emotional impact, cinematography, pacing—and different viewers weigh these aspects differently. Moreover, consumers who send recommendations often have little incentive to consider how their like will be interpreted, while the receiver may depend on it to avoid poor matches. And even a thoughtful sender cannot convey all relevant nuances through a coarse signal. Yet, recommendation systems are pervasive and people routinely rely on them. This raises fundamental questions: When are such systems useful? What shapes the value of the recommendations they generate for consumers? And how might the design incentives of platforms diverge from the interests of consumers?
The streaming example highlights three core features of recommendations that motivate our modeling approach. First, products and preferences are heterogeneous: Products come in different versions, and consumers value them differently. Second, stakes are asymmetric: Avoiding misunderstanding is more important for the receiver than for the sender. Third, recommendations are inherently coarse: The sender cannot fully convey all relevant aspects of the product. [^1] These features arise in virtually all recommendation contexts, including personal interactions with a friend. However, they are particularly pronounced on online platforms, where recommendations are typically stripped of context, reduced to simple signals, and come from (largely) anonymous senders.
@@
Our model captures these features in a parsimonious way. An agent must decide between two products, which are two-dimensional objects differing in whether they have high or low quality in each dimension. This gives rise to four distinct product versions. There is a continuum of agent types. All types agree that a product with high quality in both dimensions is worth choosing (objectively good), and that one with low quality in both is not (objectively bad)—we refer to such products as uncontroversial. By contrast, controversial products have high quality in one dimension and low quality in the other. An agent prefers the controversial product that scores highly in the dimension he cares more about (subjectively good) to the one that scores highly in the less important dimension (subjectively bad).
@@
We assume that products are experience goods so that, before consumption, each agent only knows the probability of each product version. However, he obtains a recommendation from a previous consumer of one of the products. The recommendation reveals whether the sender had a consumption experience that was sufficiently good (a buy recommendation), with payoffs above an exogenous threshold level $R$, or not (a don’t-buy recommendation). Therefore, the sender (she) always gives a buy recommendation for a good product and a don’t-buy recommendation for a bad product, independently of the threshold $R$. In addition, she sends a buy recommendation for a controversial product if her preferences for the product are strong enough.
This approach combines the features of recommendations discussed above in a stylized way: In line with her low stakes in the issue, the sender only provides a non-strategic and coarse description of her own consumption experience. [^2] In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions. However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information. By contrast, the receiver (he) of the recommendation, with his high stakes in the issue, will carefully evaluate the informational content of the recommendation using Bayesian inference, taking into account the distribution of product qualities, his own preferences, and those of the population. Further, our decision to focus only on buy and don’t-buy recommendations is a particularly simple way of capturing the ordinal nature of most (if not all) recommendation systems, which typically ask for giving “likes” or “dislikes”, rankings from 1 to 5 stars, and so forth. [^3] We do consider an extension to more than two types of recommendations in Section 6.
Recommendations contain objective and subjective content. Intuitively, from a buy recommendation a receiver learns that the product is not objectively bad and the chances that it is objectively good have increased. All receivers agree that this information makes it more attractive to buy the product. Moreover, the buy recommendation potentially influences the relative probabilities of controversial products, making it more likely that the product is the one that most consumers prefer. For consumers whose tastes are aligned with the majority of the population, this subjective content reinforces the positive objective content conveyed by the message. Thus, such receivers will definitely accept the recommendation. Contrarians, however, who understand that their tastes are not aligned with those of the rest of the population may decide not to accept the recommendation if the negative subjective content dominates the positive objective content.
Building on these simple insights into receiver behavior, we characterize both the value created by the recommendation system and the resulting demand for the recommended product. A recommendation system influences two groups of buyers: those who follow a buy recommendation (i.e., types who accept recommendations) and those who buy despite receiving a don’t-buy recommendation (i.e., types who reject recommendations). Each of these two groups may benefit from the recommendation system. For types whose preferences are sufficiently aligned with the general population, accepting buy recommendations provides value by successfully guiding them toward a product that outperforms the outside option in expectation. By contrast, for types whose preferences are misaligned with the general population, don’t-buy recommendations can provide a valuable positive signal about the fit of the product with their subjective preference.
We then turn to the question of optimal design. In practice, platforms may exert influence over how users give recommendations. For example, Uber employs a five-star system but informs riders that “5 stars means the ride went without issues,” while “4 stars or less” prompt requests for feedback or suggested improvements (Uber, 2024). [^4] In the case of Airbnb, a service provider notes that “A 5-star rating on Airbnb means that everything was as expected and that guests are happy with their stay,” while “4 stars is also considered a positive rating, but it’s not enough,” and that “3 stars or less are considered the negative feedback” (Hospitable, 2024). This suggest that platforms do not merely choose a recommendation format but also influence how users map experiences into ratings. In our framework, this mapping is governed by the threshold $R$: by shaping what counts as a positive signal, the platform effectively controls how selectively buy recommendations are given.
It is natural to assume that the platform’s choice of $R$ is guided by profit-maximization motives. As a result, the platform may aim to maximize demand, either because it receives fixed transaction fees or because higher traffic yields greater advertising revenue. We investigate the resulting choice of $R$ and compare it to what would be optimal from a consumer perspective.
We begin with a setting in which preference heterogeneity is purely idiosyncratic noise, so that the population is symmetric with respect to their valuations for the controversial products. In this case, there is no systematic preference bias in the population, so recommendations have no subjective content. We find that the value for consumers generated by the recommendation system is monotone in the threshold $R$. If good products are more likely than bad ones, the value increases with $R$, and the optimal threshold induces consumers to issue buy recommendations only when the product is objectively good. If bad products are more likely than good ones, the value decreases with $R$, and, from a consumer perspective, buy recommendations should be given unless the product is objectively bad. Either way, the designer optimally renders recommendations essentially objective in content.
In contrast with the value of the recommendation system for consumers, the demand for the recommended product is always decreasing in the threshold. Thus, a platform that seeks to maximize sales favors a low threshold, so that buy recommendations are issued except for objectively bad products. As a result, the objectives of maximizing value and maximizing demand are aligned in some environments but stand in sharp conflict in others.
We conclude the analysis of a symmetric population by examining how both the value of the recommendation system and the demand it induces depend on the economic environment. We find nuanced effects of changes in the prevalence of controversial products and in the degree of polarization in the population.
While the symmetric benchmark is analytically clean and provides useful intuition, it leaves no room for subjective content of recommendations. We therefore extend the analysis to asymmetric populations, where the controversial product versions are not valued equally on average. In this setting, value maximization no longer necessarily requires an extreme threshold which conveys the message whether a product is objectively good or objectively bad. Instead, it may be desirable to use interior thresholds which can effectively communicate whether a product is in line with majority tastes.
Next, we turn to demand maximization in the asymmetric setting. Using a combination of theoretical and numerical analysis, we find that a demand-maximizing platform usually prefers a low threshold. While this behavior may coincide with value maximization in some environments, the two objectives generally diverge in the presence of asymmetries.
Finally, various extensions of the model lead to further insights. First, we allow the sender and receiver populations to differ, as in cases where recommendations come from, say, early adopters. Our main results remain robust, and we discuss how the designer may want to shape the sender distribution. Second, we consider richer recommendation systems with more than two levels. While this added granularity might seem beneficial, we find that extreme thresholds remain optimal for value maximization in symmetric populations, and that additional levels do not help a demand-maximizing platform either. Third, we analyze the case where receivers observe multiple independent recommendations from a symmetric population. Perhaps surprisingly, access to more signals does not systematically increase value or demand: receiver behavior with more than one signal often mirrors that under a well-calibrated single-recommendation system.
Research on recommendations in the presence of preference heterogeneity is surprisingly scarce. An exception is Acemoglu et al. (2022), which argues that consumers with strong initial preferences for a product are more likely to purchase it so that the selection of recommenders is biased. Receivers should recognize this bias rather than accepting recommendations at face value. Contrary to Acemoglu et al. (2022), we focus on the determinants of the value and induced demand of recommendation systems and the design of those systems rather than the possibility and speed of learning.
@@
More broadly related, various papers deal with public goods problems in the provision of recommendations, which lead to insufficient exploration and reviewing, in particular for lesser-known niche products (e.g., Kremer et al., 2014, and Che and Hörner, 2018). [^5] Expanding on these papers, Vellodi (2022) considers the impact on market structure, showing how unbiased reviews might disadvantage entrants due to the “cold start” problem, thereby reinforcing the monopoly position of incumbents. These studies also investigate how appropriate information management can mitigate these public goods problems.
@@
While truthful recommendations are plausible in our setting, there are contexts where recommenders, firms, and platforms may have strategic incentives to bias recommendations. [^6] Bolton et al. (2013) have shown that eBay users often refrained from leaving negative feedback to avoid retaliation. Building on such concerns, Tadelis (2016) discusses how these strategic incentives shaped eBay’s platform design—for instance, the decision to disable seller evaluations of buyers—and explores broader implications for the credibility of recommendation and reputation systems. Like us, Chakraborty and Harbaugh (2010) assume that the sender possesses multidimensional information. They show how this helps her influence the behavior of the receiver in spite of a common bias. Bourreau and Gaudin (2022) consider incentives for firms to manipulate recommendations in a setting with heterogeneous products. They focus on how a platform wants to use recommendations to shift consumers to more desirable suppliers and how this affects pricing. Abstracting from product heterogeneity, Peitz and Sobolev (forthcoming) consider profit-maximizing firms that engage in sophisticated strategies to inflate their ratings. [^7] Johnen and Ng (2024) analyze a related phenomenon in a model without preference heterogeneity. They find that firms initially lower prices to attract positive ratings, later leveraging these ratings to justify price increases—effectively reducing the reliability of ratings as a quality indicator. This leads to a reduction in welfare due to inefficient product–consumer matches. The rating inflation is somewhat akin to our result that a low recommendation threshold can maximize demand and thus profits. However, in our model, this need not come at the expense of consumer welfare: when products and preferences are heterogeneous, the objectives of value and demand maximization can align, making low thresholds potentially optimal from both perspectives.
Section 2 introduces the model. In Section 3, we describe the conditions under which different receiver types accept recommendations and use this to characterize the value and demand induced by a recommendation system, respectively. Then, we in turn consider symmetric (Section 4) and asymmetric (Section 5) populations to study the optimal design of recommendation systems, both in terms of creating value and inducing demand. Section 6 discusses our results and considers extensions. Section 7 concludes. All proofs are relegated to the appendix.